Calculate the pH of a solution when 100.0 mL of a 0.100 M solution of acetic acid (CH3COOH), which has a Ka value of 1.8 × 10–5, is titrated with 110 mLs of a 0.10 M NaOH solution.
Answer #1

millimoles of CH3COOH = 100 x 0.1 = 10.0

millimoles of NaOH added = 110 x 0.10 = 11.0

11 - 10 = 1.0 millimoles exess NaOH left

[NaOH] = 1.0 / 210 = 0.00476 M

as NaOH is strong base

[NaOH] = [OH-] = 0.00476 M

pOH = - log [OH-]

pOH = - log [0.00476]

pOH = 2.32

pH = 14 - 2.32

pH = 11.68